OpenStudy (anonymous):

Please help! I'm stumped on these. The square root symbol is over the whole fraction just so you know. Simplify the following. 1. √ 4/52 2.√ 8/3n^3

OpenStudy (espex):

Recall that you can rewrite \(\sqrt{\frac{4}{52}}\) as \(\frac{\sqrt{4}}{\sqrt{52}}\)